# X-MoGen: Unified Motion Generation across Humans and Animals

### v1

@@ 1 Introduction
Motion generation aims to synthesize high-quality, physically plausible 3D motion sequences. It has become an active research area with wide-ranging applications in film production, video games, virtual/augmented reality, and robotics (Zhu et al. 2023; Sahili, Neji, and Tabia 2025). Various control modalities have been explored for motion generation, including text (Guo et al. 2022), music (Siyao et al. 2022; Tseng, Castellon, and Liu 2023; Li et al. 2024b; Shah et al. 2025; Gong et al. 2023), sketch-based input (Wang et al. 2025a; Zhong et al. 2025), and multimodal signals (Guo et al. 2025; Zhang et al. 2024b, 2025). Of these modalities, text has emerged as a particularly popular interface due to its intuitive and expressive nature. Despite recent advances, most text-driven motion generation methods focus exclusively on either human (Guo et al. 2022; Hong et al. 2025) or animal (Wang et al. 2025b) motion, treating them as independent tasks. However, unified modeling offers two key advantages (Yang et al. 2024a): (1) Unified representation learning, where aligning motion semantics across species leads to more structured and expressive latent spaces; and (2) Cross-species generalization, as diverse body morphologies and motion patterns improve transferability to unseen species and novel actions.
Previous works typically struggle to achieve unified modeling due to incompatible skeletal structures across datasets, necessitating separate models for each dataset (Guo et al. 2024; Hong et al. 2025). To address this, we introduce UniMo4D, a large-scale dataset covering humans and 114 animal species. UniMo4D standardizes skeletal topology while retaining species-specific morphology, facilitating learning in a shared representation space. This provides a foundation for joint human-animal motion modeling. However, unifying such diverse species introduces new challenges, as variations in morphology and motion dynamics make it difficult to generate consistent and realistic cross-species motions with a single model.
To address these challenges, we propose X-MoGen, a unified framework designed for human-animal motion generation. As illustrated in Figure 1, X-MoGen synthesizes motions for both humans and animals from a single text description, providing flexible and generalizable control. It also supports applications such as cross-species transformation by producing smooth and semantically consistent transitions, such as transforming a tiger into a human. X-MoGen operates in two stages. The first stage uses a conditional graph variational autoencoder to generate species-specific canonical T-pose priors, while a motion autoencoder encodes motion sequences into a compact latent space regularized by a morphological loss. In the second stage, motion is synthesized using a masked modeling mechanism (Li et al. 2023, 2024c) conditioned on the input text. To maintain structural coherence, a morphological consistency module predicts skeletal lengths from the latent representation and incorporates them into the loss computation.
@@ 2.1 Text-Driven Human Motion Generation
Text-driven motion generation methods can be broadly categorized into quantization-based and non-quantization-based approaches. Quantization-based methods (Jiang et al. 2023; Zhang et al. 2023a; Guo et al. 2024; Zhong et al. 2023; Pinyoanuntapong et al. 2024b, a; Wang et al. 2024; Liao et al. 2025; Wu et al. 2025; Lu et al. 2025) discretize motion sequences using various quantization schemes and predict motion tokens from text. However, this discretization often leads to information loss. Alternatively, non-quantization-based approaches operate directly in the continuous motion space (Zhang et al. 2023b, c, 2024a; Chen et al. 2023; Dai et al. 2024; Zhong et al. 2024; Huang et al. 2024; Hua et al. 2025; Hong et al. 2025). For example, MDM (Tevet et al. 2023) employs a diffusion model on raw motion sequences for synthesis. Subsequent works leverage latent diffusion models (Rombach et al. 2022) to improve generation quality while reducing computational cost. MARDM (Meng et al. 2025b) introduces a masked autoregressive framework for high-fidelity motion generation. However, such models are typically trained on skeletons with fixed bone lengths. The recent ShapeMove (Liao et al. 2025) supports variable bone lengths by learning additional shape features, allowing skeletal morphology to be controlled by text. In contrast, X-MoGen learns species-specific motions without relying on additional shape features, directly adapting to skeletal variations present in the data.
@@ 2.2 Text-Driven Animal Motion Generation
Generating animal motions from text is inherently more complex than human motion generation, due to the vast diversity in animal morphology and behavior. OmniMotionGPT (Yang et al. 2024b) models both human and animal motions, enabling animal synthesis via transfer learning from human motion–text pairs. Crucially, human motion serves as a key condition, guiding animal motion generation within the same category. Motion Avatar (Zhang et al. 2024c) employs large language models as high-level planners to generate 3D animal avatar meshes. AniMo (Wang et al. 2025b) introduces spatiotemporal modeling and species-aware modulation for text-driven animal motion generation. However, these methods do not explicitly address morphological variations across species. In contrast, our approach incorporates T-pose encoding and morphology-aware losses to ensure that synthesized motions accurately reflect species-specific structures.
@@ 2.3 Unified Human-Animal Motion Modeling
Building a unified framework for both human and animal motion remains challenging. WalkTheDog (Li et al. 2024a) learns a shared phase manifold for unsupervised semantic and temporal alignment across morphologically different subjects, enabling applications such as motion transfer without requiring explicit skeleton correspondences. AnyTop (Gat et al. 2025) generates natural motions aligned with a character’s biomechanical characteristics by conditioning on its skeletal structure. However, the generated motions are sampled from the training distribution and remain relatively uncontrollable. UniMoGen (Khani et al. 2025) synthesizes motions based on the input character’s style, trajectory, and motion history. In contrast to methods requiring user-specified skeletons, our X-MoGen enables cross-species motion generation solely from free-form text.
@@ 3.1 Stage 1: Feature Modeling
We propose a Conditional Graph Variational Autoencoder (CGAE) to model species-specific skeletal morphology and generalize to unseen species. CGAE integrates a conditional variational framework (Sohn, Lee, and Yan 2015) with graph convolutional encoders and decoders (Kipf and Welling 2016), explicitly treating the skeletal structure as a graph. CGAE is trained on bone-length vectors $\mathbf{b}\in\mathbb{R}^{N_{b}}$ extracted from T-poses using Bone-Length Extraction. It learns a latent distribution conditioned on species-level semantic embeddings $\mathbf{c}$ derived from BioCLIP-2’s text encoder (Gu et al. 2025). Both encoder and decoder operate on a fixed-topology graph, with bone lengths as node features.
The training objective balances reconstruction accuracy and latent regularization using the evidence lower bound (ELBO) (Kingma, Welling et al. 2013):
@@
Once trained, CGAE can generate canonical T-poses for arbitrary quadruped species. Given a condition vector $\mathbf{c}$ and a latent sample $\mathbf{z}_{\text{pose}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the decoder predicts a bone-length vector $\hat{\mathbf{b}}$. Final joint positions $\mathbf{P}$ are recovered via Forward Kinematics based on a predefined kinematic tree (Parent 2012).
@@
To encode motion sequences into an efficient and structured representation (Rombach et al. 2022), we employ an Autoencoder (AE) (Bank, Koenigstein, and Giryes 2023) that operates on raw motion inputs $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}]$. The encoder maps $\mathbf{X}$ to a temporally downsampled latent sequence $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{T}]\in\mathbb{R}^{T\times d}$, where $T=\lfloor L/4\rfloor$ and $d$ denotes the latent dimension. The encoder consists of convolutional layers with temporal strides and residual connections (He et al. 2016), while the decoder uses transposed convolutions to reconstruct the original sequence.
@@ 3.2 Stage 2: Masked Motion Generation
Recent studies have demonstrated the effectiveness of masked modeling (Li et al. 2023; Guo et al. 2024; Li et al. 2024c; Meng et al. 2025b). Building on this, we adopt a Masked Transformer (M-Trans) with a diffusion head to synthesize motion sequences. The model learns to reconstruct masked motion tokens through a denoising process, guided by text description.
@@
Textual guidance is extracted from CLIP’s text encoder (Radford et al. 2021), which provides a sentence-level feature $\mathbf{s}$ and word-level features $\mathbf{W}$. These features are projected to obtain keys and values for cross-attention:
@@
The Transformer fuses motion and text features (Vaswani et al. 2017), producing a contextualized representation $\mathbf{H}\in\mathbb{R}^{(T+1)\times d}$:
@@
The contextualized representation $\mathbf{H}$ from the M-Trans backbone serves as the conditioning signal for a diffusion-based completion head. This head is realized by a set of Multi-Layer Perceptrons (MLPs), which function as the velocity network $\mathbf{v}_{\theta}$. For each masked position $i\in\mathcal{M}$, its corresponding contextualized embedding $\mathbf{h}_{i}$ is extracted from $\mathbf{H}$. The velocity network is optimized to predict the deterministic velocity field that maps noise $\hat{\mathbf{z}}_{i}$ to the ground-truth data $\mathbf{z}_{i}$, following the flow-matching paradigm introduced by SiT (Ma et al. 2024). The optimization uses the following loss, conditioned on $\mathbf{h}_{i}$:
@@
To encourage realistic and consistent skeletal structures in generated motions, we design the Morphological Consistency Module (MCM), denoted as $f_{\text{MCM}}$. This module serves as a morphology-aware regularizer that maps the latent motion sequence $\hat{\mathbf{Z}}$ to a static vector representing the skeleton’s bone lengths. Specifically, a Gated Recurrent Unit (GRU) (Chung et al. 2014) encodes the temporal dynamics of $\hat{\mathbf{Z}}$, and its final hidden state is processed by a MLP to predict the bone-length vector. The MCM is pre-trained on the latent space of AE and remains frozen during motion generation. To ensure morphological consistency, we introduce a guidance loss that penalizes deviations from a canonical bone-length vector $\mathbf{b}$, computed from the ground-truth bone lengths:
@@ 3.3 Inference
To improve text alignment, classifier-free guidance (Ho and Salimans 2022) is applied at each denoising step to compute the guided velocity:
@@ 4 Experiments
To support joint modeling of human and animal motions, we construct UniMo4D, a unified dataset that integrates HumanML3D (Guo et al. 2022), KIT-ML (Plappert, Mandery, and Asfour 2016), and AniMo4D (Wang et al. 2025b). Among various motion representations (Guo et al. 2022; Lu et al. 2023; Meng et al. 2025b, a; Xiao et al. 2025), we adopt the essential feature groups (Meng et al. 2025b), a compact and effective scheme. Each frame is represented as a vector of dimension $N_{j}\times 3+1$, where $N_{j}$ denotes the number of joints. The root joint features include rotation velocity, linear velocity, and height (y-axis), while the remaining joints are represented by 3D positions. To create a unified representation, we first standardize the skeletal structures from all datasets. Specifically, we remap joints from KIT-ML and reorder those from AniMo4D to align with the HumanML3D convention (Petrovich, Black, and Varol 2023; Chen and UniMocap 2023), scaling bone lengths accordingly. All skeletons are then standardized to a common 25-joint topology, which incorporates three virtual tail joints to accommodate diverse morphologies. For tailless species (e.g., humans), these virtual joints are initialized to be co-located with the pelvic root. This neutral configuration ensures a consistent data structure and prevents motion artifacts, without compromising the fidelity of core body movements (Yang et al. 2023). As a result, every motion sequence in UniMo4D is represented by a skeleton with $N_{j}=25$ joints and a consistent topology.
@@
A text-motion matching model is retrained on the UniMo4D dataset to serve as the evaluation backbone (Guo et al. 2022; Meng et al. 2025b). Based on its learned feature space, motion quality, semantic alignment, diversity, and morphological consistency are evaluated. Frechet Inception Distance (FID) assesses motion realism by comparing the feature distributions of generated and real motions. R-Precision measures text-motion alignment through retrieval accuracy, while Multimodal Distance (MM-Dist) quantifies the average distance between motion and text embeddings. Diversity measures variability across motions generated from different prompts.
@@
To the best of our knowledge, X-MoGen is the first unified framework designed for cross-species motion generation involving both humans and animals. As no existing method is specifically designed for this setting, we compare our approach against recent state-of-the-art models developed for either human or animal motion generation. Specifically, we evaluate several human motion baselines, including MDM (Tevet et al. 2023), T2M-GPT (Zhang et al. 2023a), AttT2M (Zhong et al. 2023), MMM (Pinyoanuntapong et al. 2024b), MoMask (Guo et al. 2024), and MARDM (Meng et al. 2025b), along with AniMo (Wang et al. 2025b), which focuses on animal motion. All baseline models are retrained from scratch on the UniMo4D dataset using their official implementations and original data formats (Xiao et al. 2025; Meng et al. 2025a). Although these methods use different internal data formats, we extract essential motion features from their outputs to enable a fair and consistent evaluation (Meng et al. 2025b).
@@ 4.1 Quantitative Results
We evaluate X-MoGen on the UniMo4D dataset. All experiments are repeated 10 times, and we report mean values with 95% confidence intervals (Guo et al. 2024; Wang et al. 2025b). Table 1 shows the results on the test set. X-MoGen outperforms all baselines with significantly lower FID and higher R-Precision, highlighting the superior motion quality and text alignment of our method. Table 2 reports performance on the unseen species test set, which includes species excluded during training and presents a more challenging generalization scenario. X-MoGen maintains strong performance across Top-1, Top-2, FID, and MME, which are the primary metrics for evaluating motion plausibility and morphological consistency. Table 3 compares several representative motion compression methods for reconstruction. Among quantization-based approaches, VQ (Van Den Oord, Vinyals et al. 2017), FSQ (Mentzer et al. 2023), and RVQ (Lee et al. 2022; Zeghidour et al. 2021) all struggle to capture the complexity of the UniMo4D dataset, although FSQ and RVQ offer moderate improvements over vanilla VQ. Our AE does not use quantization and achieves much better reconstruction accuracy. This shows that continuous latent modeling works better for diverse motion data.

### v2

@@ 1 Introduction
Motion generation aims to synthesize high-quality, physically plausible 3D motion sequences. It has become an active research area with wide-ranging applications in film production, video games, virtual/augmented reality, and robotics (zhu2023human; sahili2025text). Various control modalities have been explored for motion generation, including text (t2m), music (siyao2022bailando), sketch-based input (zhong2025sketch2anim), and multimodal signals (guo2025motionlab; zhang2024large). Of these modalities, text has emerged as a particularly popular interface due to its intuitive and expressive nature. Despite recent advances, most text-driven motion generation methods focus exclusively on either human (t2m; hong2025salad) or animal (wang2025animo) motion, treating them as independent tasks. However, unified modeling offers two key advantages (yang2024x): (1) Unified representation learning, where aligning motion semantics across species leads to more structured and expressive latent spaces; and (2) Cross-species generalization, as diverse body morphologies and motion patterns improve transferability to unseen species and novel actions.
Previous works typically struggle to achieve unified modeling due to incompatible skeletal structures across datasets, necessitating separate models for each dataset (guo2024momask; hong2025salad). To address this, we introduce UniMo4D, a large-scale dataset covering humans and 114 animal species. UniMo4D standardizes skeletal topology while retaining species-specific morphology, facilitating learning in a shared representation space. This provides a foundation for joint human-animal motion modeling. However, unifying such diverse species introduces new challenges, as variations in morphology and motion dynamics make it difficult to generate consistent and realistic cross-species motions with a single model.
To address these challenges, we propose X-MoGen, a unified framework designed for human-animal motion generation. As illustrated in Figure 1, X-MoGen synthesizes motions for both humans and animals from a single text description, providing flexible and generalizable control. It also supports applications such as cross-species transformation by producing smooth and semantically consistent transitions, such as transforming a tiger into a human. X-MoGen operates in two stages. The first stage uses a conditional graph variational autoencoder to generate species-specific canonical T-pose priors, while a motion autoencoder encodes motion sequences into a compact latent space regularized by a morphological loss. In the second stage, motion is synthesized using a masked modeling mechanism (li2023mage; li2024autoregressive) conditioned on the input text. To maintain structural coherence, a morphological consistency module predicts skeletal lengths from the latent representation and incorporates them into the loss computation.
@@ 2.1 Text-Driven Human Motion Generation
Text-driven motion generation methods can be broadly categorized into quantization-based and non-quantization-based approaches. Quantization-based methods (jiang2023motiongpt; zhang2023generating; guo2024momask; zhong2023attt2m; pinyoanuntapong2024mmm; liao2025shape) discretize motion sequences using various quantization schemes and predict motion tokens from text. However, this discretization often leads to information loss. Alternatively, non-quantization-based approaches operate directly in the continuous motion space (rombach2022high; tevet2023human; zhang2023finemogen; zhang2024motiondiffuse; meng2025rethinking; hong2025salad). For example, MDM (tevet2023human) employs a diffusion model on raw motion sequences for synthesis. Subsequent works leverage latent diffusion models (rombach2022high) to improve generation quality while reducing computational cost. MARDM (meng2025rethinking) introduces a masked autoregressive framework for high-fidelity motion generation. However, such models are typically trained on skeletons with fixed bone lengths. The recent ShapeMove (liao2025shape) supports variable bone lengths by learning additional shape features, allowing skeletal morphology to be controlled by text. In contrast, X-MoGen learns species-specific motions without relying on additional shape features, directly adapting to skeletal variations present in the data.
@@ 2.2 Text-Driven Animal Motion Generation
Generating animal motions from text is inherently more complex than human motion generation, due to the vast diversity in animal morphology and behavior. OmniMotionGPT (yang2024omnimotiongpt) models both human and animal motions, enabling animal synthesis via transfer learning from human motion–text pairs. Crucially, human motion serves as a key condition, guiding animal motion generation within the same category. Motion Avatar (zhang2024motion) employs large language models as high-level planners to generate 3D animal avatar meshes. AniMo (wang2025animo) introduces spatiotemporal modeling and species-aware modulation for text-driven animal motion generation. However, these methods do not explicitly address morphological variations across species. In contrast, our approach incorporates T-pose encoding and morphology-aware losses to ensure that synthesized motions accurately reflect species-specific structures.
@@ 2.3 Unified Human-Animal Motion Modeling
Building a unified framework for both human and animal motion remains challenging. WalkTheDog (li2024walkthedog) learns a shared phase manifold for unsupervised semantic and temporal alignment across morphologically different subjects, enabling applications such as motion transfer without requiring explicit skeleton correspondences. AnyTop (gat2025anytop) generates natural motions aligned with a character’s biomechanical characteristics by conditioning on its skeletal structure. However, the generated motions are sampled from the training distribution and remain relatively uncontrollable. UniMoGen (khani2025unimogen) synthesizes motions based on the input character’s style, trajectory, and motion history. In contrast to methods requiring user-specified skeletons, our X-MoGen enables cross-species motion generation solely from free-form text.
@@ 3.1 Stage 1: Feature Modeling
We propose a Conditional Graph Variational Autoencoder (CGAE) to model species-specific skeletal morphology and generalize to unseen species. CGAE integrates a conditional variational framework (sohn2015learning) with graph convolutional encoders and decoders (kipf2016semi), explicitly treating the skeletal structure as a graph. CGAE is trained on bone-length vectors $\mathbf{b}\in\mathbb{R}^{N_{b}}$ extracted from T-poses using Bone-Length Extraction. It learns a latent distribution conditioned on species-level semantic embeddings $\mathbf{c}$ derived from BioCLIP-2’s text encoder (gu2025bioclip). Both encoder and decoder operate on a fixed-topology graph, with bone lengths as node features.
The training objective balances reconstruction accuracy and latent regularization using the evidence lower bound (ELBO) (kingma2013auto):
@@
Once trained, CGAE can generate canonical T-poses for arbitrary quadruped species. Given a condition vector $\mathbf{c}$ and a latent sample $\mathbf{z}_{\text{pose}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the decoder predicts a bone-length vector $\hat{\mathbf{b}}$. Final joint positions $\mathbf{P}$ are recovered via Forward Kinematics based on a predefined kinematic tree (parent2012computer).
@@
To encode motion sequences into an efficient and structured representation (rombach2022high), we employ an Autoencoder (AE) (bank2023autoencoders) that operates on raw motion inputs $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}]$. The encoder maps $\mathbf{X}$ to a temporally downsampled latent sequence $\mathbf{Z}=[\mathbf{z}_{1},\ldots,\mathbf{z}_{T}]\in\mathbb{R}^{T\times d}$, where $T=\lfloor L/4\rfloor$ and $d$ denotes the latent dimension. The encoder consists of convolutional layers with temporal strides and residual connections (he2016deep), while the decoder uses transposed convolutions to reconstruct the original sequence.
@@ 3.2 Stage 2: Masked Motion Generation
Recent studies have demonstrated the effectiveness of masked modeling (li2023mage; guo2024momask; li2024autoregressive; meng2025rethinking). Building on this, we adopt a Masked Transformer (M-Trans) with a diffusion head to synthesize motion sequences. The model learns to reconstruct masked motion tokens through a denoising process, guided by text description.
@@
Textual guidance is extracted from CLIP’s text encoder (radford2021learning), which provides a sentence-level feature $\mathbf{s}$ and word-level features $\mathbf{W}$. These features are projected to obtain keys and values for cross-attention:
@@
The Transformer fuses motion and text features (vaswani2017attention), producing a contextualized representation $\mathbf{H}\in\mathbb{R}^{(T+1)\times d}$:
@@
The contextualized representation $\mathbf{H}$ from the M-Trans backbone serves as the conditioning signal for a diffusion-based completion head. This head is realized by a set of Multi-Layer Perceptrons (MLPs), which function as the velocity network $\mathbf{v}_{\theta}$. For each masked position $i\in\mathcal{M}$, its corresponding contextualized embedding $\mathbf{h}_{i}$ is extracted from $\mathbf{H}$. The velocity network is optimized to predict the deterministic velocity field that maps noise $\hat{\mathbf{z}}_{i}$ to the ground-truth data $\mathbf{z}_{i}$, following the flow-matching paradigm introduced by SiT (ma2024sit). The optimization uses the following loss, conditioned on $\mathbf{h}_{i}$:
@@
To encourage realistic and consistent skeletal structures in generated motions, we design the Morphological Consistency Module (MCM), denoted as $f_{\text{MCM}}$. This module serves as a morphology-aware regularizer that maps the latent motion sequence $\hat{\mathbf{Z}}$ to a static vector representing the skeleton’s bone lengths. Specifically, a Gated Recurrent Unit (GRU) (chung2014empirical) encodes the temporal dynamics of $\hat{\mathbf{Z}}$, and its final hidden state is processed by a MLP to predict the bone-length vector. The MCM is pre-trained on the latent space of AE and remains frozen during motion generation. To ensure morphological consistency, we introduce a guidance loss that penalizes deviations from a canonical bone-length vector $\mathbf{b}$, computed from the ground-truth bone lengths:
@@ 3.3 Inference
To improve text alignment, classifier-free guidance (ho2022classifier) is applied at each denoising step to compute the guided velocity:
@@ 4 Experiments
To support joint modeling of human and animal motions, we construct UniMo4D, a unified dataset that integrates HumanML3D (t2m), KIT-ML (Plappert2016), and AniMo4D (wang2025animo). Among various motion representations (t2m; lu2023humantomato; meng2025rethinking; meng2025absolute; xiao2025motionstreamer), we adopt the essential feature groups (meng2025rethinking), a compact and effective scheme. Each frame is represented as a vector of dimension $N_{j}\times 3+1$, where $N_{j}$ denotes the number of joints. The root joint features include rotation velocity, linear velocity, and height (y-axis), while the remaining joints are represented by 3D positions. To create a unified representation, we first standardize the skeletal structures from all datasets. Specifically, we remap joints from KIT-ML and reorder those from AniMo4D to align with the HumanML3D convention (petrovich23tmr; chen2023unimocap), scaling bone lengths accordingly. All skeletons are then standardized to a common 25-joint topology, which incorporates three virtual tail joints to accommodate diverse morphologies. For tailless species (e.g., humans), these virtual joints are initialized to be co-located with the pelvic root. This neutral configuration ensures a consistent data structure and prevents motion artifacts, without compromising the fidelity of core body movements (yang2023unifiedgesture). As a result, every motion sequence in UniMo4D is represented by a skeleton with $N_{j}=25$ joints and a consistent topology.
@@
A text-motion matching model is retrained on the UniMo4D dataset to serve as the evaluation backbone (t2m; meng2025rethinking). Based on its learned feature space, motion quality, semantic alignment, diversity, and morphological consistency are evaluated. Frechet Inception Distance (FID) assesses motion realism by comparing the feature distributions of generated and real motions. R-Precision measures text-motion alignment through retrieval accuracy, while Multimodal Distance (MM-Dist) quantifies the average distance between motion and text embeddings. Diversity measures variability across motions generated from different prompts.
@@
To the best of our knowledge, X-MoGen is the first unified framework designed for cross-species motion generation involving both humans and animals. As no existing method is specifically designed for this setting, we compare our approach against recent state-of-the-art models developed for either human or animal motion generation. Specifically, we evaluate several human motion baselines, including MDM (tevet2023human), T2M-GPT (zhang2023generating), AttT2M (zhong2023attt2m), MMM (pinyoanuntapong2024mmm), MoMask (guo2024momask), and MARDM (meng2025rethinking), along with AniMo (wang2025animo), which focuses on animal motion. All baseline models are retrained from scratch on the UniMo4D dataset using their official implementations and original data formats (xiao2025motionstreamer; meng2025absolute). Although these methods use different internal data formats, we extract essential motion features from their outputs to enable a fair and consistent evaluation (meng2025rethinking).
@@ 4.1 Quantitative Results
We evaluate X-MoGen on the UniMo4D dataset. All experiments are repeated 10 times, and we report mean values with 95% confidence intervals (guo2024momask; wang2025animo). Table 1 shows the results on the test set. X-MoGen outperforms all baselines with significantly lower FID and higher R-Precision, highlighting the superior motion quality and text alignment of our method. Table 2 reports performance on the unseen species test set, which includes species excluded during training and presents a more challenging generalization scenario. X-MoGen maintains strong performance across Top-1, Top-2, FID, and MME, which are the primary metrics for evaluating motion plausibility and morphological consistency. Table 3 compares several representative motion compression methods for reconstruction. Among quantization-based approaches, VQ (van2017neural), FSQ (mentzer2023finite), and RVQ (lee2022autoregressive; zeghidour2021soundstream) all struggle to capture the complexity of the UniMo4D dataset, although FSQ and RVQ offer moderate improvements over vanilla VQ. Our AE does not use quantization and achieves much better reconstruction accuracy. This shows that continuous latent modeling works better for diverse motion data.